A random sample of 24 customers buying petrol was selected. From this sample the 95% confidence...

  1. A random sample of 24 customers buying petrol was selected. From this sample the 95% confidence interval estimate for the mean amount of petrol purchased per customer for the city was calculated to be between 55 and 78 litres. Determine the sample standard deviation.

    1. there is insufficient information to do this calculation

    2. 22.5

    3. 27.23

    4. 28.7
Answer #1

Ans:

sample size,n=24

df=24-1=23

critical t value=tinv(0.05,23)=2.069

Margin of error=(78-55)/2=11.5

11.5=2.069*(s/sqrt(24))

s=11.5*sqrt(24)/2.069

s=27.23